# Is Deep Learning finally better than Decision Trees on Tabular Data?

### v2

@@ E.3 Datasets with 5,000 to 10,000 instances
E.4 Datasets with 10,000 to 50,000 instances
Table 29: Classifier Performance for Instance Range: 10000-50000 and Feature Range: <=100
Dataset Phishing Adult BankMkt Elec JM1 JngChess Letter PenDigits
AutoGluon 0.997572 0.931792 0.941273 0.987260 0.770272 0.999278 0.999934 0.999725
CARTE 0.994582 0.902677 0.924664 0.909407 0.728512 0.973383 0.999440 0.999468
CatBoost 0.996482 0.930747 0.938831 0.980993 0.756611 0.976349 0.999854 0.999752
FT-Transformer 0.996760 0.914869 0.938198 0.963076 0.709321 0.999975 0.999919 0.999703
MLP 0.996991 0.928689 0.937054 0.969201 0.715620 0.999965 0.999894 0.999705
RealMLP 0.997208 0.923327 0.937031 0.961467 0.713988 0.999774 0.999914 0.999659
ResNet 0.996975 0.913790 0.935740 0.960658 0.720444 0.999956 0.999926 0.999638
SAINT 0.996746 0.920246 0.936560 0.967012 0.719464 0.999926 0.999853 0.999782
TabNet 0.996196 0.882450 0.887319 0.938656 0.674043 0.991981 0.999606 0.999753
XGBoost 0.997425 0.930482 0.938384 0.987790 0.759652 0.974087 0.999819 0.999703
XTab 0.996896 - - 0.966899 0.727984 0.999950 0.999859 0.999751
@@ E.4 Datasets with 10,000 to 50,000 instances
E.5 Datasets with more than 50,000 instances
Table 32: Classifier Performance for Instance Range: >50000 and Feature Range: <=100
@@ E.5 Datasets with more than 50,000 instances
E.6 Performance Profiles on Small and Large Data Domain
For a more fine-granular analysis of the models’ performance profiles, we conducted the analysis proposed in Section 5.6 in the small- and large-data regime separately.
Small-Data Domain.  In Figure 31, the performance profiles are shown w.r.t. the measured inference time (left) and the measured total time (right) in the small-data regime. The models CatBoost and AutoGluon yield the best performance-time ratios, with SAINT from the transformer-base models being a competitor in increasing the performance ratio $\tau$. The models FT-Transformer, ResNet, and TabNet yield similar results, where the first performs slightly better for small performance ratios, i.e., the models yield a better performance-cost ratio for a larger amount of datasets. The worst trade-off is given by TB-BERTa, where the inference time outweighs the performance.
On the right side, the performance plots are given w.r.t. the total time. In the small-data regime, as discussed in Section 5.6, the model TabPFN yields strong performance-cost ratios resulting in a superior performance followed by XGBoost, and the feedforward models MLP, and ResNet. Due to the larger training time, the fine-tuned models CARTE, XTab, and TP-BERTa are not competitive with models from other model families.
@@ E.6 Performance Profiles on Small and Large Data Domain
Large-Data Domain.  In Figure 32, the performance profiles are shown w.r.t. the measured inference time (left) and the measured total time(right) in the large-data regime. As discussed in Section 5, TabPFN is only applicable on small-data, hence, it is not included in the large-data analysis. Regarding the inference time, the models AutoGluon as an AutoML-driven approach and CatBoost from the GDBTs are superior compared to other approaches. FT-Transformer show strong results on about half the datasets used in our analysis, but cannot hold up the performance overall the whole set. The models ResNet, FT-Transformer, CARTE, and SAINT show slightly better trade-off values compared to other competitors for an increase performance ratio $\tau$. As before, TP-Berta struggles to be competitive and shows the worst performance-cost ratios.
When considering the total amount of time, the models AutoGluon (HPO) and XGBoost show the strongest performance-cost trade-offs. It is followed by CatBoost from the GDBTs family, followed by the lightweight feedforward networks, ResNet and MLP. From the fine-tuned models, XTab beats CARTE, whereas FT-Transformer wins over SAINT and Tabnet from the transformer-based approaches. Like before, TP-BERTa could not be competitive to any of the other approaches.
@@
Appendix F Ablating the Choice of Refitting
In this ablation, we explore whether refitting the model on the combined training and validation sets (after hyperparameter optimization, HPO) provides any measurable benefit. The standard procedure, as described in 3.2, uses a 10-fold nested cross-validation: we split the data into 10 folds, use 9 folds for inner cross-validation and HPO, then identify the best hyperparameter configuration and refit the model on all 9 folds before testing on the remaining fold.
The results of this ablation for CatBoost and FT-Transformer are presented below, comparing the outcomes with and without refitting.
Figure 33: Distribution of ranks for CatBoost Left and FT-Transformer Right, with and without refitting. Lower ranks indicate better performance. The spread shows the variability in rankings across datasets.
Figure 33 presents boxplots for both considered methods: CatBoost (left) and FT-Transformer (right). In the left plot, the absence of interquartile ranges indicates that CatBoost without refitting exhibits a highly consistent rank distribution, with a median rank of 1. In contrast, CatBoost with refitting has a median rank of 2. The right plot reveals a slightly different trend for the FT-Transformer. While the refitted FT-Transformer also achieves a median rank of 1, its interquartile range extends up to rank 2, indicating a broader spread in rank distribution. Meanwhile, the FT-Transformer without refitting maintains a median rank of 2. Note that one dataset was excluded for the FT-Transformer without refitting due to memory constraints.
Figure 34 illustrates the performance difference between CatBoost with and without refitting across all datasets. The results clearly show that, with only a few exceptions, CatBoost with refitting consistently outperforms its non-refitted counterpart. Similarly, Figure 35 presents the performance difference for FT-Transformer with and without refitting. Unlike CatBoost, a greater number of datasets favor the non-refitted FT-Transformer. However, overall, the majority of datasets still show improved performance with refitting.
Furthermore, we conducted a Wilcoxon signed-rank test to compare the performance of refitting versus no-refitting across multiple datasets for both CatBoost and FT-Transformer. The statistical results are summarized in Table 33.
For CatBoost, we observed an average performance improvement of $0.0079$ when refitting, with a median difference of $0.0016$. The Wilcoxon test yielded a test statistic of $180.0000$ and a highly significant p-value of $1.2985\cdot 10^{-9}$. This strongly suggests that refitting leads to a statistically significant and consistent improvement in CatBoost’s performance across datasets. Given the very low p-value $p<0.001$, we can confidently reject the null hypothesis that refitting has no effect.
Table 33: Statistical Comparison of Refit vs. No-Refit Methods
Method Pair #Datasets Avg. Diff Median Diff Wilcoxon Stat p-value
CatBoost vs. CatBoost_noRefit 68 0.0079 0.0016 180.0000 1.298511e-09
FT vs. FT_noRefit 67 0.0035 0.0004 843.0000 9.356765e-02
Table 34: Average test ROC-AUC per dataset for FT-Transformer using default refitting vs. no refitting across CV folds.
Dataset FT-Transformer FT-Transformer_norefit adult 0.914869 0.915875 analcatdata_authorship 0.999985 0.999566 analcatdata_dmft 0.576947 0.579169 balance-scale 0.999735 0.995086 bank-marketing 0.938198 0.937470 banknote-authentication 1.000000 1.000000 Bioresponse 0.820159 - blood-transfusion-service-center 0.745975 0.748119 breast-w 0.989503 0.989074 car 0.999751 0.999969 churn 0.914596 0.915300 climate-model-simulation-crashes 0.934671 0.933561 cmc 0.739402 0.736959 cnae-9 0.994497 0.994377 connect-4 0.901170 0.921978 credit-approval 0.935798 0.944236 credit-g 0.783048 0.777810 cylinder-bands 0.915494 0.826412 diabetes 0.831108 0.823379 dna 0.990937 0.989937 dresses-sales 0.620033 0.610016 electricity 0.963076 0.957884 eucalyptus 0.923933 0.911772 first-order-theorem-proving 0.796707 0.785106 GesturePhaseSegmentationProcessed 0.895166 0.799810 har 0.999685 0.999706 ilpd 0.751488 0.737753 Internet-Advertisements 0.974513 0.985391 isolet 0.998817 0.999282 jm1 0.709321 0.725904 jungle_chess_2pcs_raw_endgame_complete 0.999975 0.999861 kc1 0.783519 0.803310 kc2 0.832014 0.837281 kr-vs-kp 0.999777 0.999173 letter 0.999919 0.999886 madelon 0.747391 0.793476 mfeat-factors 0.999015 0.998560 mfeat-fourier 0.984511 0.982372 mfeat-karhunen 0.998682 0.997649 mfeat-morphological 0.970198 0.967869 mfeat-pixel 0.997451 0.998448 mfeat-zernike 0.983479 0.981858 MiceProtein 0.999973 1.000000 nomao 0.990908 0.992552 numerai28.6 0.530315 0.527963 optdigits 0.999616 0.999487 ozone-level-8hr 0.919484 0.919689 pc1 0.917591 0.840223 pc3 0.828743 0.835171 pc4 0.934944 0.944674 pendigits 0.999703 0.999668 PhishingWebsites 0.996760 0.996105 phoneme 0.965071 0.957862 qsar-biodeg 0.919584 0.914716 satimage 0.993516 0.992003 segment 0.994124 0.993598 semeion 0.995548 0.996208 sick 0.997937 0.997762 spambase 0.985969 0.983881 splice 0.992276 0.995195 steel-plates-fault 0.959182 0.962215 texture 0.999983 0.999973 tic-tac-toe 0.996152 0.996209 vehicle 0.963362 0.940233 vowel 0.999713 0.999198 wall-robot-navigation 0.999900 0.999870 wdbc 0.993967 0.986203 wilt 0.993047 0.992642
Table 35: Average test ROC-AUC per dataset for CatBoost using default refitting vs. no refitting across CV folds.
Dataset CatBoost CatBoost_norefit adult 0.930747 0.924052 analcatdata_authorship 0.999662 0.999470 analcatdata_dmft 0.579136 0.547691 balance-scale 0.972625 0.962132 bank-marketing 0.938831 0.937464 banknote-authentication 0.999935 0.999979 Bioresponse 0.885502 0.872449 blood-transfusion-service-center 0.754965 0.749848 breast-w 0.989162 0.992507 car 1.000000 0.998453 churn 0.922968 0.916146 climate-model-simulation-crashes 0.951480 0.944551 cmc 0.740149 0.735398 cnae-9 0.996316 0.994599 connect-4 0.921050 0.913372 credit-approval 0.934006 0.940661 credit-g 0.801762 0.773381 cylinder-bands 0.912070 0.867995 diabetes 0.837869 0.822365 dna 0.995028 0.994658 dresses-sales 0.595731 0.605008 electricity 0.980993 0.937421 eucalyptus 0.923334 0.916719 first-order-theorem-proving 0.831775 0.811589 GesturePhaseSegmentationProcessed 0.916674 0.779683 har 0.999941 0.999887 ilpd 0.744702 0.731536 Internet-Advertisements 0.979120 0.972513 isolet 0.999389 0.999282 jm1 0.756611 0.742362 jungle_chess_2pcs_raw_endgame_complete 0.976349 0.973983 kc1 0.825443 0.814042 kc2 0.846802 0.841593 kr-vs-kp 0.999392 0.999419 letter 0.999854 0.999802 madelon 0.937562 0.929178 mfeat-factors 0.998910 0.997917 mfeat-fourier 0.984714 0.984229 mfeat-karhunen 0.999264 0.998802 mfeat-morphological 0.965406 0.965867 mfeat-pixel 0.999422 0.999183 mfeat-zernike 0.977986 0.977831 MiceProtein 1.000000 0.999991 nomao 0.996439 0.995329 numerai28.6 0.529404 0.529350 optdigits 0.999844 0.999780 ozone-level-8hr 0.929094 0.923125 pc1 0.875471 0.850199 pc3 0.851122 0.833527 pc4 0.953309 0.945471 pendigits 0.999752 0.999728 PhishingWebsites 0.996482 0.995649 phoneme 0.968024 0.958699 qsar-biodeg 0.930649 0.928167 satimage 0.991978 0.990444 segment 0.996231 0.995441 semeion 0.998687 0.997784 sick 0.998331 0.997520 spambase 0.989935 0.988718 splice 0.995472 0.992511 steel-plates-fault 0.974350 0.968766 texture 0.999948 0.999946 tic-tac-toe 1.000000 0.999952 vehicle 0.943460 0.933394 vowel 0.999259 0.998833 wall-robot-navigation 0.999990 0.999910 wdbc 0.993813 0.991693 wilt 0.990950 0.991393
Additionally, Tables 34 and 35 present the raw results of FT-Transformer and CatBoost, respectively, in comparison to their non-refitted counterparts.

### v3

@@ G.3 Datasets with 5,000 to 10,000 instances
G.4 Datasets with 10,000 to 50,000 instances
Table 32: Classifier Performance for Instance Range: 10000-50000 and Feature Range: $\leq$ 100
Dataset Phishing Adult BankMkt Elec JM1 JngChess Letter PenDigits AutoGluon 0.997572 0.931792 0.941273 0.987260 0.770272 0.999278 0.999934 0.999725 CARTE 0.994582 0.902677 0.924664 0.909407 0.728512 0.973383 0.999440 0.999468 CatBoost 0.996482 0.930747 0.938831 0.980993 0.756611 0.976349 0.999854 0.999752 FT-Transformer 0.996760 0.914869 0.938198 0.963076 0.709321 0.999975 0.999919 0.999703 LightGBM 0.997542 0.931261 0.938470 0.989807 0.753206 0.977605 0.999825 0.999735 MLP 0.996991 0.928689 0.937054 0.969201 0.715620 0.999965 0.999894 0.999705 RealMLP 0.997208 0.923327 0.937031 0.961467 0.713988 0.999774 0.999914 0.999659 ResNet 0.996975 0.913790 0.935740 0.960658 0.720444 0.999956 0.999926 0.999638 SAINT 0.996746 0.920246 0.936560 0.967012 0.719464 0.999926 0.999853 0.999782 TabICL 0.998342 0.914430 0.940210 0.970809 0.784425 0.975471 0.999957 0.999852 TabM 0.997636 0.919662 0.941872 0.968760 0.751557 0.999985 0.999943 0.999739 TabNet 0.996196 0.882450 0.887319 0.938656 0.674043 0.991981 0.999606 0.999753 XGBoost 0.997425 0.930482 0.938384 0.987790 0.759652 0.974087 0.999819 0.999703 XTab 0.996896 - - 0.966899 0.727984 0.999950 0.999859 0.999751
@@ G.4 Datasets with 10,000 to 50,000 instances
G.5 Datasets with more than 50,000 instances
Table 35: Classifier Performance for Instance Range: >> 50000 and Feature Range: $\leq$ 100
@@ G.5 Datasets with more than 50,000 instances
Appendix H Cost vs. Efficiency Relation of Various Model Families
To observe what is the cost vs. efficiency relation of various model families, we plot the intra-search space normalized Average Distance to the Maximum (ADTM) Wistuba et al. (2016) in Figure 34, illustrating how quickly each method converges to its best solution during the HPO process.
The plot shows that XGBoost is the fastest, reaching nearly optimal performance within just 5 hours. The ResNet and MLP architecture also demonstrate notable speed, followed closely by CatBoost.
Figure 34: Intra-search space normalized average distance to the maximum over cumulative training time (seconds).
Overall, the gradient boosting methods (GBDTs) converge faster than the deep learning models. XTab, which shares the same transformer architecture as FT-Transformer, exhibits quicker convergence, likely due to its static architecture, while the FT-Transformer’s architectural components were also tuned. On the other hand, TP-BERTa is the slowest to converge, likely due to the high computational demands of its BERT-like architecture.
In Figure 35, we show the performance profiles of the models considered. We first normalize the performance values and the logarithmic time values.
Formally, let $P_{i}^{(j)}$ denote the performance of an algorithm $i$ on dataset $j$, and $T_{i}^{(j)}$ the corresponding executing time. We define $m_{\ast}^{(j)}=\max_{i}~P_{i}^{(j)}$ to be the best performance achieved across all algorithms, and $t_{\dagger}^{(j)}=\max_{i}~T_{i}^{(j)}$ as the longest runtime observed. Next, we compute for each algorithm $i$ and dataset $j$ the performance gap $\operatorname{gap}_{i}^{(j)}=(m_{\ast}^{(j)}-P_{i}(j))/m_{\ast}^{(j)}$ and the temporal gain $\operatorname{tgain}_{i}^{(j)}=(t_{\dagger}^{(j)}-T_{i}^{(j)})/t_{\dagger}^{(j)}$. Using these, we define the Performance-Time Ratio $\operatorname{ptr}_{i}^{(j)}=\operatorname{gap}_{i}^{(j)}/\operatorname{tgain}_{i}^{(j)}$ quantifying the trade-off between performance loss and time savings. To further enable comparison across datasets, we normalize the PTR values to the range $[0,1]$, such that values closer to 1 (0) indicate a better (poorer) performance-time trade-off. When computing the cumulative distribution of these normalized values, we count how many PTR values are less than or equal to a threshold $\tau\in[0,1]$ for each algorithm $i$. Hence, the resulting plot indicates how often an algorithm achieves favorable performance-time trade-offs. Curves that rise more steeply and reach higher proportions for lower $\tau$ values correspond to better overall performance-time characteristics. A red dashed corner frame in the top-left highlights the optimal trade-off.
On the left, the performance profiles are shown w.r.t. the measured inference time. The evaluation shows that feed-forward neural network models yield high performance-time ratios. TabM leads the performance-time ratios, followed by AutoGluon, RealMLP, and CatBoost. Although the AutoML framework AutoGluon shows strong performance values as discussed in more detail in Appendix F, this entails a high computational burden resulting in increased temporal costs. Both TP-Berta and TabPFN are only evaluated on a small subset of the available datasets as reflected in their low cumulative dataset coverage, where the latter shows its strong performance on the datasets it has been evaluated on by a steep increase for low values of $\tau$. The approaches SAINT, TabNet, CARTE, and XTab lie in an intermediate range, where CARTE and SAINT show slightly better performance-cost ratios compared to AutoGluon with increased values for the performance ratio $\tau$, i.e., considering a broader range of various datasets. The right plot shows the equivalent performance profiles w.r.t. the measured total time. Notably, XGBoost, AutoGluon with hyperparameter optimization, and TabM achieve strong performance-time ratios. Transformer-based models are outperformed by more lightweight models like CatBoost and ResNet which both show competitive results. From the model family encompassing foundation models, XTab shows the strongest performance. However, it is outperformed by classical GBDT approaches. TabPFN, an in-context learning model, is only applicable on small data regimes, and is therefore not competitive across the full benchmark suite.
@@
Small-Data Domain.  In Figure 36, the performance profiles are shown w.r.t. the measured inference time (left) and the measured total time (right) in the small-data regime. The models CatBoost and AutoGluon yield the best performance-time ratios, with SAINT as a transformer-based model and TabM being competitors when increasing the performance ratio $\tau$. The models FT-Transformer, ResNet, and TabNet yield similar results, where amongst these models FT-Transformer performs slightly better for small performance ratios. All of them yield a better performance-cost ratio for a larger amount of datasets. TB-BERTa exhibits the least favorable trade-off as the inference time largely outweighs the performance.
On the right side, the performance plots are given w.r.t. the total time. In the small-data regime, the model TabPFN yields strong performance-cost ratios resulting in a superior performance followed by XGBoost and the feedforward models MLP and ResNet. Due to the larger training time, the fine-tuned models CARTE, TabNet, and TP-BERTa do not match the performance ratios of models from other architectural classes.
Large-Data Domain.  In Figure 37, the performance profiles are shown w.r.t. the measured inference time (left) and the measured total time (right) in the large-data regime. As discussed in Section 5, TabPFN is only applicable to small-data, hence, it is not included in the large-data analysis. Regarding the inference time, TabM is superior to all other competitors, followed by the models AutoGluon as an AutoML-driven approach and CatBoost from the GDBTs. FT-Transformer shows strong results on about half the datasets used in our analysis, but fails to maintain performance across the entire benchmark suite and is comparable to the models TabNet, CARTE, and SAINT. This group of models shows slightly better trade-off values compared to other competitors for an increase performance ratio $\tau$. Like before, TP-Berta struggles to be competitive and shows the worst performance-cost ratios.
When considering the total amount of time, the models AutoGluon (HPO) and XGBoost show the strongest performance-cost trade-offs. It is followed by CatBoost from the GDBTs family, and the lightweight feedforward networks, ResNet and MLP. From the fine-tuned models, XTab beats CARTE, whereas FT-Transformer wins over SAINT and TabNet from the transformer-based approaches. TP-BERTa is not competitive with any of the other approaches.
Appendix I Influence of meta-feature characteristics on the predictive performance